### v2

@@ 5.3 Main Results
We benchmark several state-of-the-art methods, serving as baselines to compare with our method: 1) Caption2Music a naive baseline that employs the SpaceTimeGPT [^2] https://huggingface.co/Neleac/timesformer-gpt2-video-captioning to extract the video captions and then outputs the music by feeding captions into MusicGen (Copet et al., 2024). 2) Video2Music (Kang et al., 2023) and 3) CMT (Di et al., 2021) which botch focus on predicting MIDI notes (Rothstein, 1995) from videos while our method directly generates music signals. 4) M2UGen (Hussain et al., 2023), a strong baseline, which leverages a language model to connect vision and language, then use MusicGen (Copet et al., 2024) to generate music from language. 5) VM-NET (Hong et al., 2017), essentially different from the above methods, retrieves a piece of music from the database for a given video, while other methods predict music by training on video-music pairs.
@@
In Table 2, VidMuse, with both global and local visual modeling, exhibits impressive performance on all metrics. Specifically, compared with Video2Music (Kang et al., 2023) and CMT (Di et al., 2021), VidMuse shows the superiority in the diversity of generated music based on the density or coverage, which justifies the advantage of directly predicting music signals compared with MIDI notes. In addition, our method even outperforms the strong competitors, i.e., M2UGen. It proves that our method of predicting music directly based on video input can also achieve better performance. Furthermore, compared with a retrieval-based method i.e., VM-NET (Hong et al., 2017), VidMuse achieves a higher Imagebind score, indicating that the music generated by the learning-based strategy is more consistent with the video semantics.
5.4 User Study
In our user study, we randomly sample 600 video-music pairs from the benchmark to conduct an A/B test, which is a widely used subjective evaluation method in the music field (Donahue et al., 2023; Yuan et al., 2024), across five music generation methods: CMT, M2UGen, Caption2Music, Ground Truth (GT), and VidMuse. The test was distributed among 40 participants, ensuring each method was compared against another 60 times. The evaluation criteria are four-fold: 1) Audio quality: Refers to the sound quality of the audio track; 2) Video-music alignment: Assesses how well the music matches the visual content, e.g., a scene showing a woman crying should ideally be paired with music that sounds sad; 3) Musicality: Evaluates the aesthetic quality of the music, distinct from audio quality. For example, a piece of music may have good audio quality, but if it is out of tune, it would be considered to have poor musicality; 4) Overall assessment: Comprehensively evaluates the performance for models. Participants are asked to choose the better sample for each criterion. The user study is shown in Fig. A1, where the value at matrix[$i$][$j$] ranges from 00 to $100$, indicating the % of times listeners preferred the method in $i$-row compared to the method in $j$-column. For example, in Fig. 5 (c), the value of matrix[$2$][$4$] represents that VidMuse outperforms CMT in $77\%$ of cases in terms of Musicality. Across all criteria, our method surpasses others in more than half of the comparisons, except when compared to the ground truth. Overall, these results thoroughly validate VidMuse’s effectiveness through subjective evaluation.
5.5 Ablation Studies
In this section, we conduct comprehensive ablation studies, aiming to find the optimal design choices.
Justification of Design Choices. To validate the impact of different model design choices on our generation results and verify the effectiveness of our method, we first design two modules: a short-term modeling module (STM) and a long-term modeling module (LTM). VidMuse-STM aims to ablate the contribution of STM by removing LTM, while VidMuse-LTM utilizes only LTM. Based on the results in Table 3, we gain the insight that local information plays a more important role in the generation. By integrating global guidance with local information, we improve the alignment of the generated music with the overall video content. Second, we implement two variants with Cross-Attention with learnable Queries (CAQ) in our framework. Specifically, CAQ_SL first uses a CAQ where $K$ and $V$ are short-term features and then uses a CAQ where $K$ and $V$ are long-term features. CAQ_LS do it in an opposite order. As shown in Table 3, VidMuse outperforms two variants. It seems that the learnable query requires a deeper architecture, as demonstrated by DETR (Carion et al., 2020) and Q-Former (Li et al., 2023), but this manner reduces the model’s efficiency. Furthermore, we test a baseline using Slowfast (Feichtenhofer et al., 2019) and find its performance degenerates in Table 3. This may be because the fast path uses a higher frame rate in inputs than the slow path, but their temporal receptive fields are identical. Thus, Slowfast still lacks global guidance.
Figure 5: A/B test user study results. We introduce four criteria, i.e., audio quality, video-music alignment, musicality, and overall assessment in Sec. 5.4.
The impact of model size on performance. For the architecture of the music token decoder, we ablate three variants with different scales: small ($455$M), medium ($1.9$B), and large ($3.3$B). The medium-sized VidMuse-M is the default setting unless stated otherwise, as it achieves the best results across all metrics in Table 5. We observe the larger VidMuse-L model fails to deliver proportional improvements. This discrepancy can be partly attributed to limited GPU resources. The model’s performance is near saturation, and the larger model does not exhibit significant advantages. As a result, VidMuse-M is the optimal choice, achieving a trade-off between performance and efficiency.
Visual Encoder. We conduct ablation experiments to study the impact of various visual encoders. As shown in Table 5, we experiment with different visual encoders, including ViT (Dosovitskiy et al., 2020), CLIP (Radford et al., 2021), VideoMAE (Tong et al., 2022), and ViViT (Arnab et al., 2021). Our results show that VidMuse remains robust in processing visual information for music generation across all encoder choices. To balance computational efficiency and generation quality, we select CLIP (Radford et al., 2021) as the default visual encoder if not stated.
More ablation results are provided in Appendix A.1, where we explore the effects of pretraining data scales, different visual encoders and input settings, etc.
Table 3: Ablation studies on design choices.
@@ 5.5 Ablation Studies
  VidMuse-STM $0.898$ $45.752$ $4.915$ $1.124$ $0.470$ $0.196$
VidMuse-LTM $0.858$ $53.907$ $16.074$ $1.439$ $0.406$ $0.205$
VidMuse-CAQ_SL $0.843$ $48.940$ $3.733$ $0.947$ $0.547$ $0.163$
VidMuse-CAQ_LS $0.919$ $45.335$ $2.917$ $0.562$ $0.720$ $0.181$
VidMuse-Slowfast $1.511$ $84.683$ $10.029$ $0.266$ $0.285$ $0.118$
VidMuse 0.738 36.171 2.369 1.175 0.710 0.207
Table 4: Ablation studies on model size.
 Models Params Metrics
@@
  VidMuse-S $455$M $0.854$ $4.695$ $1.349$
VidMuse-M $1.9$B 0.843 2.413 1.487
VidMuse-L $3.3$B $0.873$ $4.041$ $1.330$
Table 5: Ablation studies on video encoders.
     Encoders           Metrics
@@
  ViViT $0.822$ 37.167 1.433 $451.83$
VideoMAE $0.778$ $37.900$ $1.074$ $360.99$
CLIP 0.753 $38.261$ $1.122$ 141.24
ViT $0.876$ $50.427$ $1.081$ $562.64$
5.6 Qualitative Analysis
In Fig. A2, our qualitative analysis underscores specific limitations for CMT and M2UGen. Specifically, CMT extracts low-level visual cues to generate symbolic music, i.e., MIDI notes. However, its training strategy for symbolic music generation results in discontinuities, particularly for slowly changing or static frames. In these situations, the model fails to predict symbolic music notes, leading to silence. M2UGen utilizes LLMs to fuse multimodal representation and then project LLMs’ embeddings into music via a text-to-music generation model. However, this approach relies on text embeddings as intermediaries, which causes the loss of visual information and restricts the model’s ability to detect nuanced visual variations. As a result, the music generated by this method usually showcases repetitive musical themes and suffers from a lack of diversity. The last row of Fig. A2 demonstrates that our Long-Short-Term (LST) approach is capable of generating music that is rich in diversity and semantically consistent with the video.
@@ 6 Conclusion
In this work, we build a rigorous pipeline to collect high-quality and diverse video-music pairs, curating a comprehensive dataset V2M. Then, we propose VidMuse, a simple yet effective method for video-to-music generation. Our approach utilizes a Long-Short-Term approach to capture both local and global visual clues in the video, allowing for the generation of contextually rich and musically diverse outputs. To validate our method, we benchmark a series of state-of-the-art methods as baselines to compare with VidMuse. Through comprehensive quantitative studies and qualitative analyses, our method has demonstrated its superiority over the existing methods.
@@
Limitations. Our work achieves a significant advancement in video-to-music generation, but it still has some limitations. First, the current implementation relies on the EnCodec model (Défossez et al., 2022) trained on a dataset that restricts the system’s capacity to a 32kHz audio sampling rate. This codec exhibits a noticeable reconstruction loss in the audio signal, potentially lowering the quality of the generated music. Second, training large models requires substantial computational resources. Our future work aims to overcome these limitations by integrating advanced codec technologies to enhance audio reconstruction fidelity, and exploring a more efficient method design.
Appendix A Appendix
A.1 Additional Experiments
Additional experiments focusing on pretraining effects, visual encoders, and codebook patterns are provided in the appendix. These parts provide insight into the decision-making process for selecting the experimental configurations within the VidMuse framework.
Pretraining Effect. Our ablation study on the effects of the data scale during finetuning, as detailed in Table A1, highlights a balance between data size and model performance. Despite not performing best in all the metrics, the model finetuned with 20k pair data emerges as our choice. The 20k data offers a compelling trade-off: it significantly improves performance across key metrics without requiring the extensive computational resources that larger datasets demand.
Table A1: Ablation studies on the ratio of finetuning data.
     Finetuning Data             Metrics
[…]
  0 0.712 $38.184$ $3.956$ $1.125$ $0.583$ $0.181$
10k $0.717$ $34.667$ $2.961$ $0.856$ $0.673$ $0.196$
20k $0.734$ 29.946 2.459 1.250 0.730 0.202
40k $0.776$ $41.075$ $3.557$ $1.094$ $0.726$ $0.195$
60k $0.828$ $40.160$ $2.844$ $0.977$ $0.660$ $0.192$
Exploration on model inputs. To explore the impact of different video sampling rates and the duration of video segments in the Short-Term module on performance, we conducted ablation studies on input FPS and short-term segment duration, detailed in Table A2. To intuitively assess the effectiveness of different settings, we employ an Average Rank (AR) metric. The AR metric ranks the results for a metric across all methods within the same table. The ranking result is from $1$ to $N$ (equals to the number of methods within the table), where $1$ is the best and $N$ is the worst. We eventually obtain AR results by averaging the ranking results for all metrics. Note that the AR results cannot be compared across different tables since this metric is designed to showcase the dominance for each method within one table clearly. From Table A2, we observe that increasing both FPS and duration tends to enhance model capabilities, suggesting that denser frame sampling yields a more detailed video representation, thereby improving music generation. Nevertheless, to balance computational costs and performance, we use a 30-second duration at 2 FPS as our optimal setting.
Codebook Pattern. The exploration of codebook interleaving patterns has attracted attention from researchers across several domains (Zeghidour et al., 2021; Wang et al., 2023; Copet et al., 2024; Yang et al., 2023a; Lan et al., 2023). In our ablation study focusing on the patterns, we find that while the Parallel and Vall-E (Wang et al., 2023) patterns align with the findings for text-to-music generation in MusicGen (Copet et al., 2024), the flattened codebook pattern does not consistently exceed the performance of the delay pattern in tasks of generating music from video. The delay pattern, notable for its relatively low computational cost, is therefore selected for our implementation. The results of this study are presented in table A3.
@@ A.1 Additional Experiments
A.2 Details of Dataset Construction

### v3

@@ 4.1 Architecture of VidMuse
Music Token Decoder. We adopt an autoregressive approach to predict the music tokens $\bar{\textbf{Y}}$ conditioned on the video segment. Music token decoder is implemented by a transformer decoder with a linear classifier. We set the latent vector size of the transformer decoder to $M$, allowing it to scale up or down the model’s size. The decoder incorporates a cross-attention mechanism that receives the visual signal $\textbf{Z}\in\mathbb{R}^{N_{s}\times P\times M}$, where $N_{s}$ is the number of frames sampled in the video segment. At each time step $t$ (where $t=1,\dots,T$), the decoder predicts the logits of current token $\bar{\textbf{Y}}_{t}\in\mathbb{R}^{K\times C}$ based on previous tokens and visual context Z. Here, $K$ denotes the number of codebooks, and $C$ represents the vocabulary size.
Audio Codec. It can convert an audio segment into discretized codebooks and, conversely, decode codebooks back into audio. The size of codebooks is $K\times T$, where $T$ denotes the length of the video. Given the Audio codec $\mathcal{C}$, we denote the Encoder as $\mathcal{C}_{\text{encode}}(\cdot)$ and the Decoder as $\mathcal{C}_{\text{decode}}(\cdot)$. In training, we need to encode the ground truth audio A into discretized tokens serving as supervise signals for the next token prediction. In the inference phase, the predicted tokens will be then decoded into music signals.
4.2 Training
where $\bar{\textbf{Y}}_{k,t,c}$ is the predicted probability of class $c$ at codebook $k$ and timestep $t$. Conditioned on the video, we train the model by predicting the next token with this loss.
5 Experiments
In this section, we elaborate on the implementation details of our experiments, and conduct massive experiments to thoroughly evaluate the efficacy of our proposed method from both subjective and objective perspectives. This is expected to provide insights for video-to-music generation.
5.1 Implementation details
Since this work does not focus on audio encoding and decoding, we use Encodec [17] for 32 kHz monophonic audio as our default compression model and use the pretrained transformer model proposed in MusicGen [10]. The training stage utilizes the AdamW optimizer [44] with a batch size of 5 samples per GPU. We sample frames from a continuous 30s video segment at 2 fps for short-term modeling and uniformly sample 32 frames from the entire video for long-term modeling. The hyperparameters are set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, with a weight decay of 0.1 and gradient clipping at 1.0. A cosine learning rate schedule is employed, incorporating a warm-up phase of 4,000 steps and an exponential moving average decay of 0.99. We use 64 H800 GPUs in pretraining and train models with $56K$ steps, which takes about 50 hours. For the finetuning stage, we utilize 32 H800 GPUs and train models with $8K$ steps, which takes about 8 hours. A top-k strategy is applied for sampling, retaining the top $250$ tokens with a temperature setting of $1.0$. In the inference stage, we set the sliding window size as $30$s, and the window’s overlap as $0.5$s.
5.2 Evaluation Metrics
To quantitatively evaluate the effectiveness of our model, we employ a series of metrics to assess different models in terms of quality, fidelity, and diversity of the generated music. These metrics include the Frechet Audio Distance (FAD), Frechet Distance (FD), Kullback-Leibler Divergence (KL), as well as Density and Coverage [55]. Additionally, we utilize the ImageBind Score [22] to examine the alignment between the video and the generated music. We acknowledge that ImageBind has limitations as it is not specifically trained on music data, but it currently seems to be a possible option for evaluating the semantic alignment between video and generated music. To fairly compare with non-public baseline models [80, 87, 65], we use evaluation metrics from their papers: beats coverage score (BCS), beats hit score (BHS), standard deviations of BCS (CSD) and BHS (HSD), and F1 scores of BCS and BHS.
5.3 Main Results
@@
We benchmark several state-of-the-art methods, serving as baselines to compare with our method: 1) Caption2Music, a naive baseline that employs the SpaceTimeGPT to extract the video captions and outputs the music by feeding captions into MusicGen [10]. 2) Video2Music [33] and 3) CMT [13] which both predict MIDI notes [58] from videos while our method directly generates music signals. 4) M2UGen [32], a strong baseline, which leverages a language model to connect vision and language, then use MusicGen [10] to generate music from language. 5) M2UGen*, a re-trained version of M2UGen using our dataset. 6) VM-NET [29], different from above methods, retrieves a piece of music from the database for a given video, while other methods predict music by training on video-music pairs.
@@
In Table 1, VidMuse, with both global and local visual modeling, exhibits impressive performance on all metrics. Specifically, compared with Video2Music [33] and CMT [13], VidMuse shows the superiority in the diversity of generated music based on the density or coverage. It justifies the advantage of directly predicting music signals compared with MIDI notes. Our method even outperforms the strong competitors, i.e., M2UGen. It proves that our method of directly predicting music based on video input can also achieve better performance. Furthermore, compared with a retrieval-based method, i.e., VM-NET [29], VidMuse achieves a higher Imagebind score, indicating that the music generated by the learning-based strategy is more consistent with the video semantics.
Table 2: Results across different benchmarks [80, 87, 37].
@@
                                                         AIST++ [37]
D2M-GAN [87] $92.3$- $91.7$- --
CDCD [88] $93.9$ $1.2$ $90.7$ 1.5 --
V2Meow [65] 100.0 0.0 $84.4$ $25.1$ $91.5$
VidMuse $99.97$ $0.3$ 96.5 $9.4$ 98.2
LORIS [80]
D2M-GAN [87] $95.6$ $9.4$ $88.7$ $19.0$ $93.1$
CDCD [88] $96.5$ $9.1$ $89.3$ $18.1$ $92.7$
LORIS[80] 98.6 6.1 $90.8$ $13.9$ $94.5$
VidMuse $96.3$ $10.6$ 95.6 8.9 95.9
TikTok [87]
D2M-GAN [87] $87.1$- $83.9$- --
CDCD [88] 91.8 -- $86.3$- --
VidMuse $79.8$ $20.0$ 97.3 $7.6$ $87.7$
Furthermore, we validate the generalization abilities of our proposed VidMuse on several different types of video-to-music benchmarks. To make fair comparisons and avoid data leakage, we check our dataset and remove the potential repeated data sample from our training set. As shown in Table 2, our VidMuse achieves comparable performance against other methods. This demonstrates that VidMuse does not overfit on our own benchmarks and exhibits strong generalization capabilities.
Table 3: Ablation studies on design choices.
@@
  VidMuse-STM $0.898$ $45.752$ $4.915$ $1.124$
VidMuse-LTM $0.858$ $53.907$ $16.074$ 1.439
VidMuse-CAQ_SL $0.843$ $48.940$ $3.733$ $0.947$
VidMuse-CAQ_LS $0.919$ $45.335$ $2.917$ $0.562$
VidMuse-Slowfast $1.511$ $84.683$ $10.029$ $0.266$
VidMuse 0.738 36.171 2.369 $1.175$
Table 4: Ablation studies on visual encoders.
@@
  ViViT $0.822$ 1.433 $451.83$ $1650$ ms $9.12$
VideoMAE $0.778$ $1.074$ $360.99$ $452$ ms $17.44$
CLIP 0.753 $1.122$ 141.24 341 ms 24.16
ViT $0.876$ $1.081$ $562.64$ $405$ ms $23.84$
Figure 4: A/B test results of the user study. We design four criteria in Sec. 5.4 to assess the subjective performance.
5.4 Subjective Comparisons with User Study
In the user study, we randomly sample 600 video-music pairs from the benchmark to conduct an A/B test, which is a widely used subjective evaluation method in the music field [14, 81]. This test includes CMT, M2UGen, Caption2Music, Ground Truth (GT), and VidMuse. The test was distributed among 40 participants, ensuring each method was compared against another 60 times. The evaluation criteria are four-fold: 1) Audio quality: Refers to the sound quality of the audio track; 2) Video-music alignment: Assesses how well the music matches the visual content, e.g., a scene showing a woman crying should ideally be paired with music that sounds sad; 3) Musicality: Evaluates the aesthetic quality of the music, distinct from audio quality. For example, a piece of music may have good audio quality, but if it is out of tune, it would be considered to have poor musicality; 4) Overall assessment: Comprehensively evaluates the performance for models. Participants are asked to choose the better sample for each criterion. The user study is shown in Fig. 4, where the value at matrix[$i$][$j$] ranges from 00 to $100$, indicating the % of times listeners preferred the method in $i$-row compared to the method in $j$-column. For example, in Fig. 4 (c), the value of matrix[$2$][$4$] represents that VidMuse outperforms CMT in $77\%$ of cases in terms of Musicality. Across all criteria, our method surpasses others in more than half of the comparisons, except when compared to the ground truth. Overall, these results thoroughly validate VidMuse’s effectiveness through subjective evaluation.
5.5 Ablation Studies
In this section, we conduct ablation studies, aiming to find the optimal design choices.
Justification of Design Choices. To validate the impact of different model design choices on our generation results and verify the effectiveness of our method, we first design two modules: a short-term modeling module (STM) and a long-term modeling module (LTM). VidMuse-STM aims to ablate the contribution of STM by removing LTM, while VidMuse-LTM utilizes only LTM. Based on the results in Table 3, we gain the insight that local information plays a more important role in the generation. By integrating global guidance with local information, we improve the alignment of the generated music with the video content. Second, we implement two variants with Cross-Attention with learnable Queries (CAQ) in our framework. Specifically, CAQ_SL first uses a CAQ where $K$ and $V$ are short-term features and then uses a CAQ where $K$ and $V$ are long-term features. CAQ_LS does it in the opposite order. As shown in Table 3, VidMuse outperforms two variants, demonstrating the efficacy of our manner. Furthermore, we evaluate a baseline that replaces the LSTV-Module with a SlowFast-like mechanism [18], where the slow path models appearance and the fast path captures temporal dynamics. As shown in Table 3, this modification degrades performance, likely because the fast path operates at a higher frame rate while sharing the same temporal receptive field as the slow path, resulting in insufficient global guidance.
Visual Encoder. We here study the impact of various visual encoders. As shown in Table 4, we experiment with different visual encoders, including ViT [16], CLIP [56], VideoMAE [67], and ViViT [4]. For fair comparisons, these encoders all use ViT-B as the backbone. The Latency and throughput are assessed with 30-second videos on the NVIDIA H800 GPU. Latency is measured with a batch size of 1, while throughput is measured with a batch size of 16. Our results show that VidMuse remains robust in processing visual information for music generation across all encoder choices. To balance computational efficiency and generation quality, we select CLIP [56] as the default visual encoder if not stated.
More ablations are provided in the appendix, including effects of the finetuning set, different input settings, etc.
@@ 6 Conclusion
In this work, we build a rigorous pipeline to collect high-quality and diverse video-music pairs, curating a comprehensive dataset V2M. Then, we propose VidMuse, a simple yet effective method for video-to-music generation. Our approach utilizes a Long-Short-Term approach to capture both local and global visual cues in the video, allowing for the generation of contextually rich and musically diverse outputs. To validate our method, we benchmark a series of state-of-the-art methods as baselines to compare with VidMuse. Through comprehensive quantitative studies and qualitative analyses, our method has demonstrated its superiority over the existing methods.
@@
Limitations. Our work achieves a significant advancement in video-to-music generation, but it still has some limitations. First, the current implementation relies on the EnCodec model [17], which sometimes exhibits a noticeable reconstruction loss for different genres of audio, potentially lowering the quality of the generated music. Second, our work solely focuses on video-to-music generation without exploring other input conditions. As such, our future work aims to overcome these limitations by integrating advanced codec technologies to enhance audio reconstruction fidelity, and exploring diverse controls during music generation.
7 Acknowledgment
The research was supported by NSFC (No. 62206234), the Early Career Scheme (ECS-Hong Kong University of Science and Technology 22201322), the Theme-based Research Scheme (T45-205/21-N), and the General Research Fund (No. 16212623) from the Hong Kong RGC, as well as the Generative AI Research and Development Center from InnoHK. We also thank Fortuna for discussions and Kyle for assisting in setting up the user study UI.
Supplementary Material
Figure A1: Distribution of music genres in the dataset, showcasing the diverse representation of genres such as electronic, classical, and jazz.
Figure A2: Distribution of instruments in the dataset, emphasizing the frequent usage of synthesizers, pianos, and drums, while also including diverse instruments such as violins and saxophones.
Figure A3: Word cloud of mood labels in the dataset, highlighting the diversity of emotions such as inspiring, happy, powerful, and dark.
8 Additional Experiments
@@
Additional experiments focusing on model inputs, codebook patterns, and finetuning effects are provided in the appendix. These parts provide insight into the decision-making process for selecting the experimental configurations within the VidMuse framework.
Exploration on model inputs. To explore the impact of different video sampling rates and the duration of video segments in the Short-Term module on performance, we conducted ablation studies on input FPS and short-term segment duration, detailed in Table A1. To intuitively assess the effectiveness of different settings, we employ an Average Rank (AR) metric. The AR metric ranks the results for a metric across all methods within the same table. The ranking result is from $1$ to $N$ (equals to the number of methods within the table), where $1$ is the best and $N$ is the worst. We eventually obtain AR results by averaging the ranking results for all metrics. Note that the AR results cannot be compared across different tables since this metric is designed to showcase the dominance of each method within one table clearly. From Table A1, we observe that increasing both FPS and duration tends to enhance model capabilities, suggesting that denser frame sampling yields a more detailed video representation, thereby improving music generation. Nevertheless, to balance computational costs and performance, we use a 30-second duration at 2 FPS as our optimal setting.
@@
Table A3: Ablation studies on the ratio of finetuning data.
     Finetuning Data             Metrics
[…]
  0 0.712 $38.184$ $3.956$ $1.125$ $0.583$ $0.181$
10k $0.717$ $34.667$ $2.961$ $0.856$ $0.673$ $0.196$
20k $0.734$ 29.946 2.459 1.250 0.730 0.202
40k $0.776$ $41.075$ $3.557$ $1.094$ $0.726$ $0.195$
60k $0.828$ $40.160$ $2.844$ $0.977$ $0.660$ $0.192$
Codebook Pattern. The exploration of codebook interleaving patterns has attracted attention from researchers across several domains [83, 71, 10, 78, 35]. In our ablation study focusing on the patterns, we find that while the Parallel and Vall-E [71] patterns align with the findings for text-to-music generation in MusicGen [10], the flattened codebook pattern does not consistently exceed the performance of the delay pattern in tasks of generating music from video. The delay pattern, notable for its relatively low computational cost, is therefore selected for our implementation. The results of this study are presented in Tab. A2.
Finetuning Effect. Our ablation study on the effects of the data scale during finetuning, as detailed in Table A3, highlights a balance between data size and model performance. Despite not performing best in all the metrics, the model finetuned with 20k pair data emerges as our choice. The 20k data offers a compelling trade-off: it significantly improves performance across key metrics without requiring the extensive computational resources that larger datasets demand. The results also validate the effectiveness of our ranking strategy based on ImageBind-AV scores (detailed in Appendix 9), showing that prioritizing videos with higher audio-visual alignment improves finetuning data quality and enhances model performance.
9 Details of Dataset Construction